OpenStudy (anonymous):

can someone explain how to do this .. the discriminant for the quadratic 0=2x^2-5x+3 is/ are _______ real solutions..

OpenStudy (anonymous):

In the quadratic equation:\[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] the discriminant is labeled by the: \[b^2-4ac\] portion

OpenStudy (anonymous):

You have a quadratic expression in the form: \[ax^2+bx+c\] Where b is negative for your specific case.

OpenStudy (haseeb96):

discriminant= b^2 -4ac from the equation a=2 ,b=-5 and c=3 now just put it in the formula and get the answer

OpenStudy (anonymous):

omg thank you guyss.... i never thought it was this easy

OpenStudy (anonymous):

My pleasure and the easy stuff is usually where people are less confidence, so be confident lalafeyza!